what is the answer to 3-4q=10q+10?
Found 2 solutions by RAY100, thrasherlax74:
Answer by RAY100(1637)

3-4q=10q+10,,,,,,,add 4q,,,,subst 10,,,,,both sides
.
-7 = 14q,,,,,divide by 14
.
- 1/2 = q
.
.
check,,,,3-4(-1/2) = 10(-1/2) +10 = 5,,,,,ok
.

Answer by thrasherlax74(11)

combine like terms. get your variables to one side and your constants to the other. answer: x=(-1/2)